# More than 200 Globular Clusters in the Milky Way, and still no super-Solar metallicity ones

### v1

@@ 2 The selected bulge Globular Cluster sample
As mentioned previously, many new GCs are recently discovered and photometrically analysed both in NIR and optical wavelengths in the MW bulge and disk. We have catalogued these objects in Table 2, where we listed their main physical parameters: positions, mean cluster proper motions (PM), heliocentric and Galactocentric distances, metallicities, ages, extinction and integrated luminosities. We also list the references from which we have taken all the information. The majority of the clusters are located inside the Galactic bulge at $R_{G}\lesssim 3$ kpc; other clusters, namely Garro 01, Riddle 15 and Gaia 2, are situated in the disk at $R_{G}>11$ kpc. Notably, clusters, located in the innermost regions, are very obscured and heavy embedded by the dust, where the differential reddening and extinction ($1.00.1<A_{Ks}<$ 1.0 mag, equivalent to $9.10.9<A_{V}<$ 9.1 mag - see Table 2) can play an important role both for the detection and the characterization of these new objects. Moreover, Figure 1 displays the position of the GCs analysed in Galactocentric coordinates, and we highlight the metallicity (changing the colour of the points) in order to show that both metal-poor and metal-rich GC populations are present inside the bulge (see also Section 4). The spatial distribution of the new sample seems to be asymmetric, preferentially towards $X>0$ kpc. This is due to the fact that most clusters present in our compilation were discovered in the VVV-VVVX footprints, but this does not mean that on the opposite side ($X<0$ kpc), other small GC candidates could live. Likely, they have not been detected yet.
@@
Figure 1: Galactocentric distribution (XY on the left, and XZ on the right) of the GCs analysed in the present work (coloured points), highlighting clusters in the inner regions of the bulge. The known Galactic GCs positions are drawn in grey points. X measured the system’s ”east-west” location, Y measured its ”north-south” location, and Z measured its distance above or below the galactic plane. The marker’s colour changes depending on the cluster’s metallicity, as indicated by the colour bar.
The black cross represents the position of the Galactic centre, while the blue star is the position of the Sun.
@@
It is important to mention some important limitations of our work. First, most of the clusters in this new catalogue have been only identified photometrically as GC candidates, therefore they deserve a detailed spectroscopic analysis for deriving robust [Fe/H] estimates as well as radial velocities and chemical abundances ([X/H]), in order to further confirm their real nature. For example, Lim et al. (2022) analysed spectroscopically 15 and 10 target stars near Camargo 1103 and Camargo 1106, respectively, two star clusters that have been recently reported as metal-poor GC candidates in the bulge (Camargo 2018). They did not find clear evidence of a grouping in radial velocity–metallicity space that would indicate the characterization of either object as metal-poor GCs. Yet, they cannot completely rule out a low probability that they only observed non-member stars. Another example may be represented by Garro 01, since its classification is ambiguous. It was classified as GC with [Fe/H] $=-0.7$ and age $\approx 11$ Gyr (as written in Table 2) by Garro et al. (2020) and recently as possible open cluster with [Fe/H] $=-0.3$ and age $=4$ Gyr by Pace et al. (2023). However, they cannot confirm the Garro 01 cluster nature either. Here, we assume that all these targets (Table 2) are genuine GCs. Second, incomplete colour-magnitude diagrams (CMDs), which do not reach the main-sequence turn-off point (MSTO), affects the age and luminosity estimates. In particular, in many cases ages are derived using the vertical method (see for example the work by Garro et al. 2022b) or identifying RR Lyrae stars belonging to the clusters, thus fixing a lower limit, but indicating large error bars (Table 2). This also implies that we cannot well-classified them as (young) GCs or old open clusters (OCs), even if it is not well-established a net separation between these two classes.
A third caveat is represented by the fact that many of these clusters, took into consideration in the present work, are located in very obscured regions, where the differential reddening, high extinction and stellar crowding may complicate the analysis, introducing uncertainties in the derivation of the main physical parameters. However, these problems are highly discussed by the authors of the papers (see Table 2 - last column).
Finally, taking into account the errors in the photometry, we can consider that this first work on the new upgraded MW GC system can give us further clues about the GC system itself, as well as indications about the MW’s history.
This study includes the information collected to date, and the sample considered here contains both the known and well-characterised GCs and those from recently studied and confirmed as high-probability GCs (Table 2). In this way, we are upgrading the MW GC system, therefore we expect that this has an impact on the MW GC luminosity function (LF - Sect. 3) and the metallicity distribution function (MD - Sect. 4).
@@
Table 1: Catalogue of the 37 recently discovered GCs in the MW. We list (where available) their main physical parameters: position, mean cluster PMs, extinction in $K_{s}$ band, distances, metallicity, age, total luminosity, and their references.
@@ 3 Milky Way globular cluster luminosity function
Until now, all the information that we have about the MW GC system were based on the known and well-characterised Galactic GCs, listed in previous compilations such as Harris (1996) and Vasiliev & Baumgardt 2021 catalogues. Discovering and identifying these clusters has been easy over time, as most of them are very bright, massive, highly concentrated and host a larger number of stars. On the other hand, new GCs (Table 2) are mainly faint, widespread, embedded in very obscured regions. Therefore, the detection of low-luminosity GCs, possible thanks to powerful surveys, is only able within our Galaxy, while they are not visible in external galaxies. Hence, this may allow us to explore a region of the LF not accessible before, going towards the tail of this feature and increasing statistics at these luminosity ranges.
@@
The MW GC LF can be used as a distance indicator (Rejkuba 2012) or to constrain the GCs’ history (Kruijssen & Zwart 2009). In our case, we employ it for the second scope. As mentioned before, many dynamical processes can destroy a GC, and all these events can have an effect on the MW GC LF, affecting in particular the faint end tail of the distribution. We investigate this in order to see how the known MW GC LF changes if we include all the recently discovered GCs.
@@
As we can see from Table 2 and Figure 3 (orange lines), the LF for the new GCs extends towards much lower luminosities. Actually, we found a double peaked distribution: the first peak at $M_{V}^{new}\leavevmode\nobreak\ =\leavevmode\nobreak\ -6.06\pm 0.80$ mag, and a second one at $M_{V}^{new}\leavevmode\nobreak\ =\leavevmode\nobreak\ -3.83\pm 0.62$ mag (this is a tentative one since the $1\sigma$ Poisson error is ±$\pm 2$ GCs, therefore the statistical significance of this peak is very low). Moreover, we constructed the LF for all known and well-characterised GCs (blue lines), using the $M_{V}$ values from the 2010 compilation of the Harris (1996) catalogue. A less pronounced double-peak is present for this sample at $M_{V}^{old}=-7.46\pm 0.20$ and $M_{V}^{old}=-4.35\pm 0.20$, in agreement with the literature (e.g., Garro et al. 2021a). At this point, if we merge the two samples, thus upgrading the ”typical” GC LF (green lines), we find that the distribution is slightly shifted towards lower luminosities, with the bright peak at $M_{V}^{up}=-7.00\pm 1.30$ mag and the faint one at $M_{V}^{up}=-4.10\pm 0.49$ mag, with an indication of elongated faint end tail. However, the significance and physical meaning of the LF peaks must be taken with caution, because it is well known that the GCLF is asymmetric when plotted in magnitude bins (see e.g. very similar histograms in Fall & Zhang 2001). In this instance, the statistical significance of the second peak is not as robust, given the limited number of clusters within that range (specifically, at $-5<M_{V}<-3$, there are 26 objects with Poisson error for each bin of ±$\pm 2-3$), however it does not disappear even when we include the Poisson error. Hence, this observation might imply the existence of numerous other faint clusters yet to be discovered, likely obscured by the considerable challenges in their detection. This study could serve as an initial step in delving into the analysis of this faint segment of the MW GC LF.
@@
As demonstrated by several works, the GC LFs is not Gaussian nor universal for all galaxies (Rejkuba 2012). The shape of the GC LF deviates from the Gaussian symmetric distribution, since it shows a longer tail toward the faint end. This trend is related to the main dynamical processes that can destroy a GC (which can be amplified especially when galaxy merger events are involved), and they are: (i) the mass loss due to stellar evolution, such as the supernova explosions (Lamers et al. 2010); (ii) dynamical friction (Alessandrini et al. 2014); (iii) tidal shock heating by passages through the bulge or disk (Piatti & Carballo-Bello 2020); and (iv) evaporation due to two-body relaxation (Madrid et al. 2017). Moreover, Kruijssen & Zwart (2009) suggested that the efficiency of the dynamical evolution are linked to the initial cluster LF and mass function, which depend on the M/L parameter differently. In turn, M/L depends on the cluster mass dependent mass-loss rate (dissolution time $t_{dis}\propto M^{0.7}$ and $t_{dis}\propto t_{rh}^{0.75}$, where $t_{rh}$ is half-mass relaxation time, Kruijssen & Zwart 2009). This indicates that if M/L increases with mass and luminosity, the low-mass and low-luminosity GCs, present in our compilation, are more strongly depleted in low-mass stars. This because the massive stars are more segregated in the central regions, whereas the low-mass stars are preferentially located in the cluster’s outskirts, where the cluster’s potential well is less strong than the central regions, thus the chance of losing stars is most probable in the external regions of the cluster. Hence, the estimation of the luminosity, which we are using (see Table 2, and reference there in) is reliable, since it mainly comes from evolved and brighter stars. Nonetheless, a dependence of M/L on mass is not obvious, since it is still not clear what the mass function can change depending on the cluster’s metallicity, as suggested by Baumgardt (2016).
@@
Another important feature that we have to include is the fact that the MW GC LF may be bimodal, even if the second peak does not contain a high number of clusters. However, we want to compare the obtained MW GC LF with the M31 galaxy’s LF (Peacock et al. 2010; Huxor et al. 2014; Garro et al. 2021a) – which shows a double peaked distribution – in order to see if we can deduce some important clues about our Galaxy. By this comparison, we can deduce that the second peak may be related to GCs, which have been accreted or survived disruption processes (Huxor et al. 2014). This may be supported by the fact that the new clusters that are very faint ($M_{V}>-5.0$), they are also younger than 11 Gyr, metal-richer than -1.0, and located in the Bulge, where dynamical processes are very strong (Figure 4). Precisely, Ferrero 54 and Gaia 2 located at larger $R_{G}$ are disk-like GCs, and Patchick 125 is the ancient clusters in our compilation. If this is true, our targets have lost a high percentage of their stars, therefore likely those low-luminosity (and low-mass) GCs may be what remains of more massive progenitors. This result is in agreement with recent studies that suggest that the Galactic stellar halo has assembled from tidally-disrupted dwarf galaxies (see e.g. Belokurov et al. 2018; Conroy et al. 2019; Kruijssen et al. 2019). However, most of the stellar streams, or the accretion events in general, have been discovered across the Galactic halo (Ibata et al. 2001; Odenkirchen et al. 2001; Mateu 2023). This may represent a ”selection effect”, given that studying the halo is easier than exploring the bulge, just because the stellar crowding and extinction are irrelevant there. Finding new objects in the bulge, which may be what remains of disrupted GCs, and why not accreted, it could pave the way for the study of ex-situ populations also in the bulge. Nonetheless, we need numerical simulations and spectroscopic analysis that can confute this hypothesis. For that, we can speculate that this new MWGC LF may be the result of different processes, such as a higher number of accreted objects, efficient destruction due to the above-mentioned dynamical processes that occurred both in the Galactic halo and in the bulge, and/or incompleteness of our catalogue since many other faint and obscured objects still remain to be identified. In particular, we suspect that with the discovery and confirmation as real GCs of other candidates, which are as well low-luminosity, maybe the LF can change once completeness is achieved.
@@
Figure 3: Luminosity functions (LFs). The left panel shows three histograms depicting the LFs for the known MW GC system by Harris (1996) (2010 edition), using the dashed blue line, for the GCs analysed in this work and listed in Table 2, using a dashed orange line and the sum of the two mentioned distributions in the green line. In the right panel, we show the LFs using the KDE technique for the three samples.
@@ 4 Metallicity distribution
The metallicity distribution (MD) is an important tool to yield information about the age and formation history of a stellar system, especially when this is investigated in different environments. As shown in Figure 5, we compared the new GC MD (green line) with the MDs of known GCs using three samples: the 2010 compilation of the Harris (1996) catalogue (blue line), the Mészáros et al. (2020) catalogue (orange line), which includes especially halo+disk GCs, and our compilation of bulge GCs listed in Table 2. The first sample contains 157 GCs and their [Fe/H] values are both spectroscopic and photometric estimates [^3] Here extensive references are given for the Harris’ catalogue, https://physics.mcmaster.ca/%7Eharris/mwgc.ref, whereas the second includes 31444Mészáros et al. (2020) excluded from their catalogue those GCs that show significant differential reddening (E(B-V) $>0.4$), which introduces errors in the estimation of atmospheric parameters and consequently in the chemical abundances estimates. known GCs and the [Fe/H] is taken from SDSS-IV APOGEE-2, while third catalogue is created using [Fe/H] and [Mg/Fe] taken from the literature, as specified in the fourth column.
@@
Mészáros et al. (2020) found that the [Fe/H] metallicities that they derived are on average 0.162 dex higher than those from the Harris catalogue, so we subtracted this quantity to their [Fe/H] in order to bring the MDs to the same scale. Due to the fact that the histograms depend on the bin size, we also preferred to construct a kernel density estimation (KDE) diagram, using the epanechnikov kernel, a smoothing of 0.25, and normalised per area. The KDE code provides us the highest peak of the metallicity distributions.Independently, in order to better identify the peaks of the distributions, we overplotted a densogram for each sample, using the acos scaling [^5] We constructed the densogram with TOPCAT, and we used the acos scaling only because the peaks of the distributions were more marked than the linear scaling. which is in agreement with the previous method. A third and more robust method that we implemented is the Gaussian mixture model (GMM). We tested different Gaussian components, and that the best fit is returned with a 2-Gaussian fit, from which we obtained the following peaks and errors. As well, these results are in agreement within the errors with the previous estimations. In all cases, we found double-peaked distributions (see Figure 5):
  • for the Mészáros et al. (2020) sample: the metal-poor peak is at [Fe/H] $=-1.49\pm 0.40$ dex and the metal-rich peak is at [Fe/H] $\leavevmode\nobreak\ =-0.56\pm 0.46$ dex. The latter is not clearly visible in the histogram distribution, due to the choice of bin size, but it is clear in the KDE diagram; finally a third peak at very metal-poor range is detectable at [Fe/H] $\leavevmode\nobreak\ =-2.20\pm 0.38$ dex;
  • for the Harris (1996) (2010 edition) sample: the metal-poor peak is at [Fe/H] $=-1.59\pm 0.36$ dex and the metal-rich peak is at [Fe/H] $\leavevmode\nobreak\ =-0.59\pm 0.35$ dex, in agreement with the literature (e.g., Minniti 1995; Carretta et al. 2009; Smith 2000; Ness & Freeman 2016; Dias et al. 2016; Garro et al. 2021a);
  • for the bulge GCs sample (see Table 2): the metal-poor peak is at [Fe/H] $=-1.03\pm 0.20$ and the metal-rich peak is at [Fe/H] $=-0.38\pm 0.20$;
  • for the new sample, containing clusters in Table 2: metal-poor peak is at [Fe/H] $=-1.38\pm 0.35$ dex and metal-rich is at [Fe/H] $\leavevmode\nobreak\ =-0.59\pm 0.23$ dex, with another tentative peak at [Fe/H] $\leavevmode\nobreak\ =-0.12\pm 0.20$ dex.
@@
Table 2: Parameters for bulge GCs. We construct our compilation of 20 bulge GCs, collecting [Fe/H] and [Mg/Fe] values from the literature.
@@
Figure 5: Metallicity distributions. On the right and in the middle panels we show the KDEs and histograms for the known MW GCs for the Mészáros et al. (2020) (orange lines), Harris (1996) (2010 edition - blue lines), bulge GCs (Table 2 - magenta lines), and our work (green lines) samples, respectively. In the right panel, we also show the densograms in order to better discriminate the higher peaks (yellower colours) of the distributions. On the left panel, we show histograms for different samples of Bulge FS, as specified by the legend. The grey areas, in the middle and left panels, are used to point out that at metallicity [Fe/H] $>0$ we do not have any GCs, while a higher percentage of Bulge FS is located at that range.
@@
Basically, the distributions, obtained using these different samples, are in agreement between them, since both metal-rich and metal-poor peaks show similar values. In this way, we can confirm the metallicity bimodality of GCs, as expected since it is well studied in the literature (e.g. Tonini 2013). However, we can use the MDs in order to (i) understand something more about our compilation; (ii) apply known models in order to draw conclusions regarding the nature of our sample.
@@
Figure 6 displays that the new GCs sample, which are mainly situated in the MW bulge, follows the same trend as the inner (bulge) GCs. In addition, we can see a farther tentative peak at [Fe/H] $\approx-0.1$. This may suggest that our sample may be contaminated by younger stellar systems, which should be more metal-rich than older population, as GCs are. This might be expected given that age estimates are approximate for all our targets, and OCs may be included in our sample. Examples could be: Ferrero 54, which was classified as MW disk GC by Garro et al. (2023), with an age $=11$ Gyr and [Fe/H] $=-0.2$; and, Kronberger 49, located at $R_{G}=1.14$ kpc, with an age $=11$ Gyr and [Fe/H] $=-0.2$. For that reason, we need both spectroscopic analysis and deeper CMDs, in order to constrain their parameters.

### v2

@@ 2 Sample of selected bulge GCs
As mentioned previously, many new GCs are recently discovered and photometrically analyzed both in NIR and optical wavelengths in the MW bulge and disk. We have cataloged these objects in Table 2, where we listed their main physical parameters: positions, mean cluster proper motions (PM), heliocentric and Galactocentric distances, metallicities, ages, extinction, and integrated luminosities. We also list the references from which we have taken all the information. The majority of the clusters are located inside the Galactic bulge at $R_{G}\lesssim 3$ kpc; other clusters, namely Garro 01, Riddle 15, and Gaia 2, are situated in the disk at $R_{G}>11$ kpc. Notably, clusters, located in the innermost regions, are very obscured and heavy embedded by the dust, where the differential reddening and extinction ($1.00.1<A_{Ks}<$ 1.0 mag, equivalent to $9.10.9<A_{V}<$ 9.1 mag - see Table 2) can play an important role both for the detection and the characterization of these new objects. Moreover, Figure 1 displays the position of the GCs analyzed in Galactocentric coordinates, and we highlight the metallicity (changing the color of the points) in order to show that both metal-poor and metal-rich GC populations are present inside the bulge (see also Section 4). The spatial distribution of the new sample seems to be asymmetric, preferentially toward $X>0$ kpc. This is due to the fact that most clusters present in our compilation were discovered in the VVV-VVVX footprints, but this does not mean that on the opposite side ($X<0$ kpc), other small GC candidates could live. Likely, they have not been detected yet.
@@
Figure 1: Galactocentric distribution (XY on the left, and XZ on the right) of the GCs analyzed in the present work (colored points), highlighting clusters in the inner regions of the bulge. The known Galactic GCs positions are drawn in grey points. X measured the system’s ”east-west” location, Y measured its ”north-south” location, and Z measured its distance above or below the galactic plane. The marker’s color changes depending on the cluster’s metallicity, as indicated by the color bar. The black cross represents the position of the Galactic centre, while the blue star indicates the position of the Sun.
@@
It is important to mention some important limitations of our work. First, most of the clusters in this new catalog have been only identified photometrically as GC candidates, therefore they deserve a detailed spectroscopic analysis for deriving robust [Fe/H] estimates as well as radial velocities and chemical abundances ([X/H]), in order to further confirm their real nature. For example, Lim et al. (2022) carried out a spectroscopic analysis of 15 and 10 target stars near Camargo 1103 and Camargo 1106, respectively, two star clusters that have been recently reported as metal-poor GC candidates in the bulge (Camargo 2018). They did not find any clear evidence of a grouping in radial velocity-metallicity space that would indicate the characterization of either object as metal-poor GCs. Yet, they cannot completely rule out a low probability that they only observed non-member stars. Another example may be represented by Garro 01, since its classification is ambiguous. It was classified as GC with [Fe/H] $=-0.7$ and age $\approx 11$ Gyr (as written in Table 2) by Garro et al. (2020) and recently as possible open cluster with [Fe/H] $=-0.3$ and age $=4$ Gyr by Pace et al. (2023). However, they cannot confirm the Garro 01 cluster nature either. Here, we assume that all these targets (Table 2) are genuine GCs. Second, incomplete color-magnitude diagrams (CMDs) that do not reach the main sequence turnoff point (MSTO), affects the age and luminosity estimates. In particular, in many cases ages are derived using the vertical method (see for example the work by Garro et al. 2022b) or identifying RR Lyrae stars belonging to the clusters, thus fixing a lower limit, but indicating large error bars (Table 2). This also implies that we cannot aptly classify them as (young) GCs or old open clusters (OCs), even if a net separation is clearly established between these two classes.
A third caveat is represented by the fact that many of these clusters (considered in the present work) are located in very obscured regions, where the differential reddening, high extinction and stellar crowding may complicate the analysis, introducing uncertainties in the derivation of the main physical parameters. However, these problems are highly discussed by the authors of the papers (see Table 2, last column). Finally, taking into account the errors in the photometry, we can consider that this first work on the newly updated MW GC system can offer further clues about the GC system itself, as well as indications about the MW’s history.
This study includes the information collected to date, and the sample considered here contains both the known and well-characterized GCs and those from recently studied and confirmed as high-probability GCs (Table 2). In this way, we are upgrading the MW GC system, therefore we expect that this has an impact on the MW GC luminosity function (LF, see Sect. 3) and the metallicity distribution function (MD, see Sect. 4).
@@
Table 1: Catalog of the 37 recently discovered GCs in the MW. We list (where available) their main physical parameters: position, mean cluster PMs, extinction in $K_{s}$ band, distances, metallicity, age, total luminosity, and their references.
@@ 3 Milky Way globular cluster luminosity function
Until now, all the information that we have about the MW GC system were based on the known and well-characterized Galactic GCs, listed in previous compilations such as the catalogs from Harris (1996) and Vasiliev & Baumgardt 2021. Discovering and identifying these clusters has been easy over time, as most of them are very bright, massive, highly concentrated, and hosting a larger number of stars. On the other hand, new GCs (Table 2) are mainly faint, widespread, embedded in very obscured regions. Therefore, the detection of low-luminosity GCs, possible thanks to powerful surveys, is only able within our Galaxy, while they are not visible in external galaxies. Hence, this may allow us to explore a region of the LF not accessible before, going toward the tail of this feature and increasing statistics at these luminosity ranges.
@@
The MW GC LF can be used as a distance indicator (Rejkuba 2012) or to constrain the GC history (Kruijssen & Zwart 2009). In our case, we employ it for the second scope. As mentioned earlier in this paper, many dynamical processes can destroy a GC, and all these events can have an effect on the MW GC LF, affecting in particular the faint end tail of the distribution. We investigate this in order to see how the known MW GC LF changes if we include all the recently discovered GCs.
@@
As we can see from Table 2 and Figure 3 (orange lines), the LF for the new GCs extends toward much lower luminosities. Actually, we found a double peaked distribution: the first peak at $M_{V}^{new}\leavevmode\nobreak\ =\leavevmode\nobreak\ -6.06\pm 0.80$ mag and a second one at $M_{V}^{new}\leavevmode\nobreak\ =\leavevmode\nobreak\ -3.83\pm 0.62$ mag (this is a tentative one since the $1\sigma$ Poisson error is ±$\pm 2$ GCs, therefore the statistical significance of this peak is very low). Moreover, we constructed the LF for all known and well-characterized GCs (blue lines), using the $M_{V}$ values from the 2010 compilation of the Harris (1996) catalog. A less pronounced double-peak is present for this sample at $M_{V}^{old}=-7.46\pm 0.20$ and $M_{V}^{old}=-4.35\pm 0.20$, in agreement with the literature (e.g., Garro et al. 2021a). At this point, if we merge the two samples, thus upgrading the ”typical” GC LF (green lines), we find that the distribution is slightly shifted toward lower luminosities, with the bright peak at $M_{V}^{up}=-7.00\pm 1.30$ mag and the faint one at $M_{V}^{up}=-4.10\pm 0.49$ mag, with an indication of elongated faint end tail. However, the significance and physical meaning of the LF peaks must be taken with caution, because it is well known that the GCLF is asymmetric when plotted in magnitude bins (see, e.g., very similar histograms in Fall & Zhang 2001). In this instance, the statistical significance of the second peak is not as robust, given the limited number of clusters within that range (specifically, at $-5<M_{V}<-3$, there are 26 objects with Poisson error for each bin of ±$\pm 2-3$); however it does not disappear even when we include the Poisson error. Hence, this observation might imply the existence of numerous other faint clusters yet to be discovered, likely obscured by the considerable challenges in their detection. This study could serve as an initial step in delving into the analysis of this faint segment of the MW GC LF.
@@
As demonstrated by several works, the GC LFs is not Gaussian nor universal for all galaxies (Rejkuba 2012). The shape of the GC LF deviates from the Gaussian symmetric distribution, since it shows a longer tail toward the faint end. This trend is related to the main dynamical processes that can destroy a GC (which can be amplified especially when galaxy merger events are involved), which are: (i) the mass loss due to stellar evolution, such as the supernova explosions (Lamers et al. 2010); (ii) dynamical friction (Alessandrini et al. 2014); (iii) tidal shock heating by passages through the bulge or disk (Piatti & Carballo-Bello 2020); and (iv) evaporation due to two-body relaxation (Madrid et al. 2017). Moreover, Kruijssen & Zwart (2009) suggested that the efficiency of the dynamical evolution are linked to the initial cluster LF and mass function, which depend on the M/L parameter differently. In turn, M/L depends on the cluster mass dependent mass-loss rate (dissolution time $t_{dis}\propto M^{0.7}$ and $t_{dis}\propto t_{rh}^{0.75}$, where $t_{rh}$ is half-mass relaxation time, Kruijssen & Zwart 2009). This indicates that if M/L increases with mass and luminosity, the low-mass, and low-luminosity GCs, present in our compilation, are more strongly depleted in low-mass stars. This because the massive stars are more segregated in the central regions, whereas the low-mass stars are preferentially located in the cluster’s outskirts, where the cluster’s potential well is less strong than the central regions; thus, the chance of losing stars is most probable in the external regions of the cluster. Hence, the estimation of the luminosity, which we are using (see Table 2, and reference there in) is reliable, since it mainly comes from evolved and brighter stars. Nonetheless, a dependence of M/L on mass is not obvious, since it is still not the impact of the mass function can change depending on the cluster’s metallicity, as suggested by Baumgardt (2016).
@@
Another important feature that we have to include is the fact that the MW GC LF may be bimodal, even if the second peak does not contain a high number of clusters. However, we want to compare the obtained MW GC LF with the M31 galaxy’s LF (Peacock et al. 2010; Huxor et al. 2014; Garro et al. 2021a) – which shows a double peaked distribution – to see if we can deduce some important clues about our Galaxy. By this comparison, we can deduce that the second peak may be related to GCs, which have been accreted or survived disruption processes (Huxor et al. 2014). This may be supported by the fact that the new clusters that are very faint ($M_{V}>-5.0$), they are also younger than 11 Gyr, more metal-rich than -1.0, and located in the bulge, where dynamical processes are very strong (Figure 4). Precisely, Ferrero 54 and Gaia 2 located at larger $R_{G}$ are disk-like GCs, and Patchick 125 is the ancient clusters in our compilation. If this is true, our targets have lost a high percentage of their stars, therefore likely those low-luminosity (and low-mass) GCs may be what remains of more massive progenitors. This result is in agreement with recent studies that suggest that the Galactic stellar halo has assembled from tidally disrupted dwarf galaxies (see e.g., Belokurov et al. 2018; Conroy et al. 2019; Kruijssen et al. 2019). However, most of the stellar streams, or the accretion events in general, have been discovered across the Galactic halo (Ibata et al. 2001; Odenkirchen et al. 2001; Mateu 2023). This may represent a ”selection effect,” given that studying the halo is easier than exploring the bulge, just because the stellar crowding and extinction are irrelevant there. Finding new objects in the bulge, which may be what remains of disrupted GCs, and considering why they were not accreted, could pave the way for the study of ex situ populations also in the bulge. Nonetheless, we need numerical simulations and spectroscopic analysis that can confute this hypothesis. For that, we can speculate that this new MWGC LF may be the result of different processes, such as a higher number of accreted objects, efficient destruction due to the above-mentioned dynamical processes that occurred both in the Galactic halo and in the bulge, and/or incompleteness of our catalog since many other faint and obscured objects still remain to be identified. In particular, we suspect that with the discovery and confirmation as real GCs of other candidates, which are also of low luminosity, it may be possible that the LF can change once completeness is achieved.
[…]
figurec
@@
Figure 3: Luminosity functions (LFs). The left panel shows three histograms depicting the LFs for the known MW GC system by Harris (1996) (2010 edition), using the dashed blue line, for the GCs analyzed in this work and listed in Table 2, using a dashed orange line and the sum of the two mentioned distributions in the green line. In the right panel, we show the LFs using the KDE technique for the three samples.
@@ 4 Metallicity distribution
The metallicity distribution (MD) is an important tool for obtaining information about the age and formation history of a stellar system, especially when this is investigated in different environments. As shown in Figure 5, we compared the new GC MD (green line) with the MDs of known GCs using three samples: the 2010 compilation of the Harris (1996) catalog (blue line), the Mészáros et al. (2020) catalog (orange line), which includes, in particular, the halo+disk GCs, and our compilation of bulge GCs listed in Table 2. The first sample contains 157 GCs and their [Fe/H] values are both spectroscopic and photometric estimates [^3] Extensive references are given for the Harris’ catalog at https://physics.mcmaster.ca/%7Eharris/mwgc.ref, whereas the second includes 31444Mészáros et al. (2020) excluded from their catalog GCs exhibiting significant differential reddening (E(B-V) $>0.4$), which introduces errors in estimations of atmospheric parameters and, consequently, the chemical abundance estimates. known GCs and the [Fe/H] is taken from SDSS-IV APOGEE-2, while the third catalog is created using [Fe/H] and [Mg/Fe] taken from the literature, as specified in the fourth column.
@@
Mészáros et al. (2020) found that the [Fe/H] metallicities that they derived are on average 0.162 dex higher than those from the Harris catalog, so we subtracted this quantity from their [Fe/H] to bring the MDs to the same scale. Due to the fact that the histograms depend on the bin size, we also preferred to construct a kernel density estimation (KDE) diagram, using the epanechnikov kernel, a smoothing of 0.25, and normalized per area. The KDE code provides us the highest peak of the metallicity distributions.Independently, to better identify the peaks of the distributions, we overplotted a densogram for each sample, using the acos scaling [^5] We constructed the densogram with TOPCAT, and we used the acos scaling only because the peaks of the distributions were more marked than the linear scaling. in agreement with the previous method. A third and more robust method that we implemented is the Gaussian mixture model (GMM). We tested different Gaussian components and that the best fit came from a two-Gaussian fit, from which we obtained the following peaks and errors. In addition, these results are in agreement within the errors with the previous estimations. In all cases, we found double-peaked distributions (see Figure 5): (i) for the Mészáros et al. (2020) sample, the metal-poor peak is at [Fe/H] $=-1.49\pm 0.40$ dex and the metal-rich peak is at [Fe/H] $\leavevmode\nobreak\ =-0.56\pm 0.46$ dex. The latter is not clearly visible in the histogram distribution, due to the choice of bin size, but it is clear in the KDE diagram. Finally, a third peak at very metal-poor range is detectable at [Fe/H] $\leavevmode\nobreak\ =-2.20\pm 0.38$ dex; (ii) for the Harris (1996) (2010 edition) sample, the metal-poor peak is at [Fe/H] $=-1.59\pm 0.36$ dex and the metal-rich peak is at [Fe/H] $\leavevmode\nobreak\ =-0.59\pm 0.35$ dex, in agreement with the literature (e.g., Minniti 1995; Carretta et al. 2009; Smith 2000; Ness & Freeman 2016; Dias et al. 2016; Garro et al. 2021a); (iii) for the bulge GCs sample (see Table 2), the metal-poor peak is at [Fe/H] $=-1.03\pm 0.20$ and the metal-rich peak is at [Fe/H] $=-0.38\pm 0.20$; (iv) for the new sample containing clusters in Table 2, metal-poor peak is at [Fe/H] $=-1.38\pm 0.35$ dex and metal-rich is at [Fe/H] $\leavevmode\nobreak\ =-0.59\pm 0.23$ dex, with another tentative peak at [Fe/H] $\leavevmode\nobreak\ =-0.12\pm 0.20$ dex.
@@
Table 2: Parameters for bulge GCs. We construct our compilation of 20 bulge GCs, collecting [Fe/H] values from the literature.
@@
Figure 5: Metallicity distributions. Right and in the middle panels we show the KDEs and histograms for the known MW GCs for the Mészáros et al. (2020) (orange lines), Harris (1996) (2010 edition - blue lines), bulge GCs (Table 2 - magenta lines), and our work (green lines) samples, respectively. In the right panel, we also show the densograms in order to better discriminate the higher peaks (yellower colors) of the distributions. On the left panel, we show histograms for different samples of bulge FSs, as specified by the legend. The grey areas, in the middle and left panels, are used to point out that at metallicity [Fe/H] $>0$ we do not have any GCs, while a higher percentage of bulge FSs is located at that range.
@@
Basically, the distributions, obtained using these different samples, are in agreement between them, since both metal-rich and metal-poor peaks show similar values. In this way, we can confirm the metallicity bimodality of GCs as expected, since it is well studied in the literature (e.g., Tonini 2013). However, we can use the MDs in order to (i) understand something more about our compilation and (ii) apply known models in order to draw conclusions regarding the nature of our sample.
@@
Figure 6 displays that the new GCs sample, which are mainly situated in the MW bulge, follows the same trend as the inner (bulge) GCs. In addition, we can see a farther tentative peak at [Fe/H] $\approx-0.1$. This may suggest that our sample may be contaminated by younger stellar systems, which should be more metal-rich than older population, as GCs are. This might be expected given that age estimates are approximate for all our targets, and OCs may be included in our sample. Examples could be: Ferrero 54, which was classified as MW disk GC by Garro et al. (2023), with an age $=11$ Gyr and [Fe/H] $=-0.2$; and Kronberger 49 located at $R_{G}=1.14$ kpc, with an age $=11$ Gyr and [Fe/H] $=-0.2$. For that reason, we need both spectroscopic analysis and deeper CMDs to constrain their parameters.